OpenStudy (anonymous):

Equations with fractions; Show work! 3/2 (x + 8) = 6x

OpenStudy (e.cociuba):

To solve for the unknown "x", let's first get rid of the parenthesis by multiplying out the 3/2 leaving us with: 3x/2 + 12 = 6x

OpenStudy (e.cociuba):

Subtract 3x/2 from both sides of the equation leaving: 12 = 6x - 3x/2 = 9x/2

OpenStudy (e.cociuba):

Multiply both sides by 2 leaving: 24 = 9x

OpenStudy (e.cociuba):

Divide both side by 9 leaving: x = 24 /9 or 8/3 or 2 and 2/3

OpenStudy (anonymous):

Can you please show me it worked out on one comment!

OpenStudy (e.cociuba):

3/2x + (3/2*8)=6x 3/2x + 12 = 6x 3/2x -6x = -12 x (3/2 - 6) = -12 x (3-12)/2 = -12 -9/2 x = -12 x = -12 * -2/9 x = 24 /9 or 8/3 or 2 and 2/3